### v1

# Consecutive Model Editing with Batch alongside HooK Layers

Abstract
As the typical retraining paradigm is unacceptably time- and resource-consuming, researchers are turning to model editing in order to seek an effective, consecutive, and batch-supportive way to edit the model behavior directly. Despite all these practical expectations, existing model editing methods fail to realize all of them. Furthermore, the memory demands for such succession-supportive model editing approaches tend to be prohibitive, frequently necessitating an external memory that grows incrementally over time. To cope with these challenges, we propose COMEBA-HK, a model editing method that is both consecutive and batch-supportive. COMEBA-HK is memory-friendly as it only needs a small amount of it to store several hook layers with updated weights. Experimental results demonstrate the superiority of our method over other batch-supportive model editing methods under both single-round and consecutive batch editing scenarios. Extensive analyses of COMEBA-HK have been conducted to verify the stability of our method over 1) the number of consecutive steps and 2) the number of editing instances. Our code will be released via https://github.com/anonymous.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) (Chung et al., 2022; OpenAI, 2023; Black et al., 2022; Touvron et al., 2023) have been demonstrated to be capable of recalling factual knowledge about the real world (Brown et al., 2020; Petroni et al., 2020). Nevertheless, researchers also reveal that LLMs often fail to recall the most up-to-date knowledge or information and some specialized knowledge if they are not periodically updated (Liska et al., 2022; Agarwal & Nenkova, 2022; Lazaridou et al., 2021). Despite the fact that fresh and customizable knowledge is highly desired in many areas, such as text generation, question-answering, reasoning, etc., updating the model via retraining is both time- and resource-consuming. Additionally, researchers have uncovered that well-trained LLMs do make mistakes. One popular sort of mistake is called hallucination (Tonmoy et al., 2024), which means that LLMs generate text based on ”hallucinated” fake knowledge. Although many researchers have tried to mitigate this issue (Qiu et al., 2023; Mündler et al., 2023; Kang et al., 2023; Varshney et al., 2023), the strategy to fix this bug remains unclear. Therefore, researchers have started to seek an efficient approach that could edit LLMs in a customizable, cost-effective way.
To this end, recent years have witnessed many efforts in investigating the model-editing techniques to bypass the retraining paradigm and edit the LLMs directly (Meng et al., 2022; Hartvigsen et al., 2022; Li et al., 2023; Mitchell et al., 2022a, b). Accordingly, several new datasets (e.g., ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023)) and corresponding metrics (e.g., reliability, generality, locality, portability (Yao et al., 2023)) are proposed to facilitate the development in this field. However, these methods either require extra expensive training of a meta-network (Mitchell et al., 2022a), or a classifier (Mitchell et al., 2022b), which causes time and resources overhead, or demands an external memory of explicit edit instances for reference (Mitchell et al., 2022b), which inevitably escalates the memory requirement. Further, most of the methods proposed so far are only targeted for single-round editing, where the model is rolled back to the initial state after each edit step. We believe this deviates from the application scenario in reality since most users anticipate an editing approach that is succession-supportive and batch-supportive, i.e., a method that advocates consecutive batch editing.
In light of these issues, we propose a novel method, named COMEBA-HK, which performs COnsecutive Model Editing with BAtch alongside HooK layers. Specifically, COMEBA-HK supports consecutive batch editing and utilizes the hook layers to separate weight change from the original model weight. COMEBA-HK does not need any training or large explicit external memory that stores editing instances. It only needs a tiny amount of memory to collect the optimized weight in the hook layer. To achieve this, we propose a new transformer memory updating mechanism that supports consecutive batch editing settings and design a simple yet effective local editing scope identification technique used in the hook layer that can accurately detect the inputs in the local editing scope. We demonstrate the effectiveness of our method via extensive experiments on ZsRE and COUNTERFACT datasets using two popular autoregressive language models, GPT2-XL and GPT-J (6B). Both the single-round batch settings and consecutive batch settings are included, with the number of consecutive batch editing instances ranging from 1k to 10k. An analysis of the editing scope identification has also been conducted to validate the method. Beyond all these, we implement comprehensive ablation studies to verify the validity of each component and discuss the optimal parameter settings in the method.
2 Preliminaries of Model Editing
As defined by Yao et al. (2023), the task of model editing is to efficiently modify an initial base model $f_{\theta}$ into an edited model $f_{\theta^{\prime}}$ whose responses to a particular set of input instances $\mathcal{X}_{t}$ are adjusted as desired without affecting the responses of the model to other instances. The intended edit descriptor is denoted as $(x_{t},y_{t})$, where $x_{t}\in\mathcal{X}_{t}$ and $f_{\theta}(x_{t})\neq y_{t}$. The post-edit model $f_{\theta^{\prime}}$ is supposed to produce the expected output to an intended edit instance $x_{t}$, while preserving the original output to other instances:
$f_{\theta^{\prime}}(x)=\left\{\begin{array}[]{l}y_{t}\quad\quad~{}~{}\text{if}~{}x\in\mathcal{X}_{t}\\ f_{\theta}(x)\quad\text{if}~{}x\notin\mathcal{X}_{t}\end{array}\right..$ (1)
In particular, there are three standard criteria for model editing, namely Reliability, Generality, and Locality (Yao et al., 2023; Mitchell et al., 2022a, b). Suppose the prediction of the original model to the statement “What is the native language of Joe Biden?” is “French”, and the expected post-edit model prediction is “English”. To verify the Reliability, we use the same original statement as input and then assess whether the post-edit model predicts “English” as desired. For Generality, a rephrased statement “What is the native language of Joe Biden?” could be inputted into the edited model to assess whether the output of the model remains as “English”. Locality suggests that the model output of an irrelevant statement like “What is the native language of Donald Trump?” should remain unaffected, which means that the post-edit model should output whatever the initial model output to this statement.
The current problem settings of model editing can be generally categorized into three groups (Yao et al., 2023):
1) Single instance Editing evaluates the post-edit model performance when only one single knowledge update is performed:
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (2)
2) Batch Editing evaluates the post-edit model performance in a more realistic scenario where multiple knowledge pieces are modified simultaneously:
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum\nolimits_{t=1}^{n}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (3)
where $n\leq\mid{\mathcal{X}_{t}}\mid$ is the batch size and it varies for different methods (Meng et al., 2023; Mitchell et al., 2022a, b; Meng et al., 2022).
3) Sequential Editing requires every single edit to be performed successively, and evaluation has to be conducted after a series of knowledge updates (Hartvigsen et al., 2022):
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum\nolimits_{t=1}^{\mid{\mathcal{X}_{t}}\mid}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (4)
In this work, we investigate a new and more practical problem setting for model editing, namely Consecutive Batch Editing, which aims at executing the editing in a consecutive and batch-edits-supportive way.
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum_{s=0}^{\lceil\mid{\mathcal{X}_{t}}\mid/n\rceil}\sum_{t=s\times n}^{\min((s+1)\times n,\mid{\mathcal{X}_{t}}\mid)}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ ,$ (5)
where $s$ represents the consecutive editing step.
3 Method
We first discuss our method under the single-layer consecutive batch editing setting. Explicitly, we first discuss the process of extending the single-layer updating mechanism in MEMIT (Meng et al., 2023) from a scenario of single-round batch editing to consecutive batch editing. Then, we describe the motivation, definition, and application of the hook layer and the local editing scope identification operation employed in the hook layer. The practicality of the operation is also clarified. Finally, the motivation and procedure for broadening the method from the single-layer case to the multi-layer case are discussed.
3.1 Single-Layer Consecutive Batch Editing
3.1.1 Batch Editing Menchanism
Meng et al. (2023) demonstrate an effective single-layer editing method using minimal squared error. Although it supports multiple edits on a single round, the updates do not account for scenarios involving consecutive updates. In this section, we extend this approach to include consecutive scenarios. Strang (2022) suggests Eq.6 can be optimized if we can find a $W_{0}$ that satisfies Eq.7, where $K_{0}=[k_{1}|k_{2}|...|k_{n}]$ and $V_{0}=[v_{1}|v_{2}|...|v_{n}]$ are set of key-value pairs we want to store into the layer.
$W_{0}=\operatorname*{argmin}_{W}\sum\nolimits_{i=1}^{n}\norm{Wk_{i}-v_{i}}^{2}$ (6)
$W_{0}K_{0}K_{0}^{T}=V_{0}K_{0}^{T}\ .$ (7)
Thanks to the well-conducted pre-training procedure for most of the available transformer-based LLMs, we can assume that the pre-training weight $W_{0}$ satisfies Eq.7, i.e., serves as the optimal solution for Eq.6.
Unlike Meng et al. (2023), we define a successive mass-editing objective:
$W_{1}=\operatorname*{argmin}_{W}(\sum_{i=1}^{r}\norm{Wk_{i}-v_{i}}^{2}+\sum_{i=r+1}^{r+u}\norm{Wk_{i}-v_{i}}^{2})$ (8)
$W_{1}[K_{1}\,K_{2}][K_{1}\,K_{2}\,]^{T}\,=\,[V_{1}\,V_{2}][K_{1}\,K_{2}]^{T}\ ,$ (9)
where $K_{1}=[k_{1}|k_{2}|\dots|k_{r}](r\geq n)$ and $V_{1}=[v_{1}|v_{2}|\dots|v_{r}]$ is the set of key-value pairs that have been updated and $K_{2}=[k_{r+1}|k_{r+2}|...|k_{r+u}]$ and $V_{2}=[v_{r+1}|v_{r+2}|...|v_{r+u}]$ is the set of key-values that are going to be edited. Therefore, the objective (Eq.8) indicates that we want an optimal $W_{1}$ that successfully updates the new associations while maintaining the old key-value pairs.
Further expanding Eq.9:
$(W_{1}+\Delta)(K_{1}K_{1}^{T}+K_{2}K_{2}^{T})=(V_{1}K_{1}^{T}+V_{2}K_{2}^{T})$ (10)
$\begin{split}W_{1}K_{1}K_{1}^{T}&+\Delta K_{1}K_{1}^{T}+W_{1}K_{2}K_{2}^{T}\\ &+\Delta K_{2}K_{2}^{T}=V_{1}K_{1}^{T}+V_{2}K_{2}^{T}\end{split}$ (11)
In a real consecutive editing scenario, $r$ increases and starts with $n$, and each batch-editing iteration is optimized through the objective (Eq.8). Hence, we can conclude that $W_{1}K_{1}K_{1}^{T}=V_{1}K_{1}^{T}$. Subtracting it from Eq.11, we get:
$\Delta K_{1}K_{1}^{T}+W_{1}K_{2}K_{2}^{T}+\Delta K_{2}K_{2}^{T}=V_{2}K_{2}^{T}$ (12)
$\Delta=RK_{2}^{T}C_{accu}^{-1}\ ,$ (13)
where $R=(V_{2}-W_{1}K_{2})$ is the residual error evaluated on the most recent updated weights (Note that $W_{1}=W_{0}$ if and only if $r=n$) and $C_{accu}=(K_{1}K_{1}^{T}+K_{2}K_{2}^{T})$ the accumulation sum of editing keys’ outer product. Note that $K_{1}K_{1}^{T}$ is computed via:
$K_{1}K_{1}^{T}=K_{0}K_{0}^{T}+K^{\prime}K^{\prime T}\ ,$ (14)
where $K_{0}$ is the set of pre-training keys that have been contained in the pre-training weight, $K^{\prime}=[k_{n+1}|k_{n+2}|...|k_{r}]$ denotes the updated keys proceeding to current editing step. We follow (Meng et al., 2023) to model $K_{0}K_{0}^{T}$ as the uncentered covariance of some randomly sampled inputs:
$K_{0}K_{0}^{T}=\lambda E[kk^{T}]\ .$ (15)
Note that the $\lambda$ represents a factor that balances the pre-trained and the whole updated associations. We follow the definitions and applications of keys and values in (Meng et al., 2023, 2022), where keys are the activations at the last token of the subject and values are gradient-descent optimized vectors that maximize the model’s prediction for the target object (Fig.1).
Figure 1: Single layer update with hook layer (residual connections are omitted).
3.1.2 Hook Layer
Yao et al. (2023) demonstrated that those editing methods that directly modify the model parameter in place struggle with sequential editing. Specifically, the locality decreases drastically when the number of iterations increases. Meanwhile, those methods that freeze the model parameters show more stable performance over iterations. This indicates that it might be helpful to separate the editing change from the model itself. However, it is too memory-costly to straightforwardly make another copy of the model or apply an external memory (Mitchell et al., 2022b; Hartvigsen et al., 2022) for a consecutive batch editing scenario. Therefore, a potentially cost-effective trade-off might be only copying part of the model layers and editing merely on those copied layer weights. This allows us to have access to both the original and edited outputs so that we can design criteria to precisely swap specific parts of the original output for the edited outputs whenever necessary.
With all these motivations, we introduce the hook layer (Fig.1), which takes the original model layer weights as the weight initialization and is responsible for all editing weight alteration in the whole editing process of COMEBA-HK. It serves a similar idea with the forward hook function defined in popular ML libraries like PyTorch, which adjusts the original forward layer output based on some predefined criterion. Theoretically, the hook layer can be hung up on any target linear layer in the transformer, Nevertheless, we mainly focus on the critical path identified in (Meng et al., 2023, 2022) as they are verified to be crucial for fact associations storage in the autoregressive language model.
As shown in Fig.1, there are generally two sorts of hook layers in this work, namely, the Temporary hook layer and the Validated hook layer. The temporary hook layer is mainly applied during the process of editing. It replaces the whole original output with the output from the hook layer so that the residual with respect to the hook layer weights instead of the original layer weights is computed. The hook layer weights are then updated (Eq.13) using the calculated residual and the accumulated sum of keys’ outer product. Validated hook layers are employed after each batch editing step at the layer, and they inherit the updated weights from the temporary hook layer. We describe the operation conducted by it in the section 3.1.3.
Figure 2: Multiple layer update with hook layer (Attention module and the first FFN are omitted).
3.1.3 Local Editing Scope Identification
Outlier Demonstration
Given the original outputs produced by the model layer weights and the edited outputs generated by hook layer weights, we need to decide when and which part of the original outputs to swap over. The ideal solution is to only switch those parts of outputs whose keys have been updated to the hook layer weights and leave other parts unchanged. To this end, we first detect the output parts that have their keys updated. Suppose $k_{i}\in K_{1},v_{i}\in V_{1}$ is an association that has been updated to $W_{1}$, and $k_{j}\notin K_{1}$ is a key that is not included in the updated associations. We show empirically in section 4.4 that $\parallel{W_{1}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{1}k_{j}-W_{0}k_{j}}\parallel$ holds. This implies that when the hook layer with updated weight $W_{h}$ receives an input $\hat{K}\in R^{m\times n}$ (batch dimension is ignored for simplicity) that contains an edited key $k_{i}\in\hat{K}\cap K_{1},\,k_{i}\in R^{m}$, then we should have $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ for all $k_{j}\in\hat{K}-\hat{K}\cap K_{1}$, which means that $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$ would be outliers among $\{\parallel{W_{h}k_{x}-W_{0}k_{x}}\parallel:\!\forall k_{x}\in\hat{K}\}$. Hence, detecting outputs of the updated keys can be transferred to detecting the outliers in the L2-norm distribution of inputs. We used the standardization to find the outliers (Fig.1), which applies the standardization technique to L2-norm vectors of inputs and determines outliers via a predefined threshold $\alpha$. Concretely, for the inputs $\hat{K}$, we first compute the L2-norm vector $M^{l}\in R^{n}$:
$P^{l}=W_{0}\hat{K}\quad O^{l}=W_{h}\hat{K}$ (16)
$M^{l}_{i}=\parallel(O^{l}_{i}-P^{l}_{i})\parallel\ .$ (17)
Then, we standardize $M^{l}$ and select the swap location by comparing it with $\alpha$.
Threshold $\alpha$ Determination
We denote $Z_{x}=\max((M^{l}-\mu)/\sigma)$ as the maximum z-score of an input $\hat{K}$. Since the $Z_{x}$ varies for different instances (section 4.4) and is likely to shift as the consecutive editing steps grow, it is not reasonable to set $\alpha$ as a fixed real number. Therefore, we determine the $\alpha$ dynamically during the editing process:
$\alpha_{s}=\begin{cases}\alpha_{z}&\text{if}\;s=1,\\ \min(\alpha_{c},\alpha_{s-1})&\text{otherwise}\end{cases}\ ,$ (18)
where $s\geq 1$. Specifically, the $\alpha$ is first initialized to a pre-selected value $\alpha_{z}$. At each consecutive editing step $s$, for the batch of inputs in this step, we calculate $Z_{x}$ for each single instance and select the minimum $Z_{x}$ in the batch (i.e., the supremum) as the candidate $\alpha_{c}$. The $\alpha_{s}$ is finally determined to be the minimum between the candidate $\alpha_{c}$ and the previous value $\alpha_{s-1}$. In practice, we set $\alpha_{z}=2.2$.
Table 1: Single round batch editing results with batch size 30.
Method Model ZsRE COUNTERFACT
Reliability Generality Locality Reliability Generality Locality
FT (Meng et al., 2022) GPT2-XL 0.1685 0.1634 0.7155 0.0027 0.0034 0.8518
MEND (Mitchell et al., 2022a) 0.0357 0.03413 0.1993 0.0013 0.0003 0.0422
SERAC (Mitchell et al., 2022b) 0.4608 0.3007 0.1962 0.0010 0.0045 0.8225
MEMIT (Meng et al., 2023) 0.6119 0.4997 0.9751 0.8101 0.2767 0.9580
COMEBA-HK 0.8221 0.6661 0.9940 0.8828 0.4038 0.9766
FT (Meng et al., 2022) GPT-J 0.2257 0.2177 0.9919 0.0037 0.0034 0.9957
MEND (Mitchell et al., 2022a) 0.1050 0.0918 0.1225 0.1728 0.0351 0.0858
SERAC (Mitchell et al., 2022b) 0.2766 0.2633 0.6052 0.0020 0.0028 0.9999
MEMIT (Meng et al., 2023) 0.9340 0.7045 0.9647 0.9957 0.4229 0.9525
COMEBA-HK 0.9759 0.7241 0.9910 0.8794 0.4276 0.9817
3.2 Multiple-layer Consecutive Batch Editing
Given the designed single-layer editing procedure, there exists a risk that the single-layer hook fails to detect the updated keys. Suppose $k_{i}$ is an updated key; failure to detect $k_{i}$ indicates that the output corresponds to $k_{i}$ at this single layer would be the same as the original output $W_{proj}k_{i}$, which consequently leads to the failure of edition to $k_{i}$. We believe one potential solution is to apply the hook to multiple model layers rather than a single model layer because the latter layer grasps the chance to capture the edited keys missed by proceeding layers. Furthermore, (Zhu et al., 2020) showcased that minimizing the magnitude of parameter change is helpful for improving the robustness of the model. Thus, we expand our work to multiple layers (Fig.2).
We first find the desired object vector $v_{i}$ following a similar procedure in (Meng et al., 2023). However, the optimization is not based on the original model, but the model hung with the validated hook that inherits the most recently updated hook weights from the previous editing step. After $v_{i}$ is found, we start to update the hook weight at each layer.
At each batch editing step, all the hook layers are initialized to temporary hook layers, which substitute the entire original output to output from hook layers. The purpose of doing this is to ensure that the residual regarding the hook layer weights rather than the original model weights are calculated. Then, the residual is distributed evenly to each layer, and the alteration $\Delta^{l}$ to the parameter at each layer is found in a layer-increasing iterative manner with keys, and residuals recomputed at each iteration (Fig.2). The reason for the recomputation of keys and residuals is that the layer-increasing alteration approach will affect the keys and residuals in the latter layer. For each layer, once the hook layer weight is updated, the hook layer is changed from a temporary hook layer to a validated hook layer to facilitate the computation of the keys and residuals in the latter layer. After the whole editing process is completed, the validated hook layers with the ultimately updated weights are hung on the model to shape the final edited model.
4 Experiments
4.1 Experiment Setups
Datasets
We use the ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023) datasets with the split provided in EasyEdit [^1] https://github.com/zjunlp/EasyEdit/tree/main for evaluation. ZsRE is a question-answering dataset that uses back-translation to generate equivalent neighborhoods. It is initially used in factual knowledge evaluation and later adopted in model editing by (Mitchell et al., 2022a). COUNTERFACT is a challenging dataset that focuses on counterfactual information that has a low prediction score compared to correct facts. It builds out-of-scope data by replacing the subject entity with a similar description that shares the same predicate. Further details are provided in Appendix.A.
Baselines
For baselines , we adopt several batch-supportive editing methods, including SERAC (Mitchell et al., 2022b), MEND (Mitchell et al., 2022a), MEMIT (Meng et al., 2023), and fine-tuning (FT) technique used in (Meng et al., 2022; Yao et al., 2023), which only fine-tune a specific layer identified by Rome (Meng et al., 2022) instead of all layers to ensure a fair comparison. We choose large autoregressive language models GPT2-XL and GPT-J (6B) as our base models. All baselines are implemented using the EasyEdit (Yao et al., 2023; Zhang et al., 2024; Wang et al., 2023; Cheng et al., 2023; Mao et al., 2023; Zhang et al., 2023) library. Further details of the datasets and implementation are given in the Appendix A.
Evaluation Metrics
We employ three popular editing evaluation metrics defined in (Yao et al., 2023; Huang et al., 2023; Cao et al., 2021), i.e., reliability, generality, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
$\mathbb{E}_{(x_{t},y_{t})\in\{(x_{t},y_{t})\}}\{\arg\max\nolimits_{y}f_{\theta^{\prime}}(y|x_{t})=y_{t}\}\ .$ (19)
The editing should also edit the equivalent neighbor of the instance $N(x_{t},y_{t})$ (e.g. rephrased descriptions). This metric is named generality and is evaluated by the average accuracy on the neighbors of the edit cases:
$\mathbb{E}_{(x_{t}^{\prime},y_{t}^{\prime})\in\{N(x_{t},y_{t})\}}\{\arg\max\nolimits_{y}f_{\theta^{\prime}}(y|x_{t}^{\prime})=y_{t}^{\prime}\}\ .$ (20)
Despite the editing, those instances that are irrelevant to the edit cases $\{O(x_{t},y_{t})\}$ should not be affected. This evaluation is called locality (also known as specificity) and is measured by the proportion of unchanged predictions between the initial model and the post-edit model:
$\mathbb{E}_{(x_{t}^{\prime},y_{t}^{\prime})\in\{O(x_{t},y_{t})\}}\{f_{\theta^{\prime}}(x_{t}^{\prime})=f_{\theta}(x_{t}^{\prime})\}\ .$ (21)
Table 2: Consecutive batch editing results with batch size 10 on 1k samples.
Method Model ZsRE COUNTERFACT
Reliability Generality Locality Reliability Generality Locality
FT (Meng et al., 2022) GPT2-XL 0.0379 0.0248 0.0660 0.0100 0.0100 0.0600
MEND (Mitchell et al., 2022a) 0.0002 0.0001 0.0012 0.0001 0.0000 0.0008
MEMIT (Meng et al., 2023) 0.3488 0.3296 0.7074 0.5600 0.3700 0.3100
COMEBA-HK 0.6691 0.5611 0.9723 0.8600 0.3800 0.5900
FT (Meng et al., 2022) GPT-J 0.2353 0.2170 0.5527 0.0200 0.0200 0.7200
MEND (Mitchell et al., 2022a) 0.0005 0.0001 0.0002 0.0100 0.0100 0.0100
MEMIT (Meng et al., 2023) 0.6336 0.4890 0.7480 0.7500 0.4500 0.4200
COMEBA-HK 0.7989 0.6129 0.9652 0.9500 0.4100 0.8000
4.2 Evaluation on Single-round Batch Editing
We first test the effectiveness of our method under basic single-round batch editing settings, i.e., the model is rolled back to the initial state after each batch editing. Both MEMIT and COMEBA-HK need to set the parameter $\lambda$, the balance factor between pre-trained and newly updated associations. According to (Meng et al., 2023), higher $\lambda$ helps preserve the original model behavior (locality) but could be harmful to reliability and generality, and the best overall performance is found at around $\lambda=10^{4}$. However, with the intent to verify whether our method comprehensively improves the editing, that is, could accept lower $\lambda$ to assign higher weight for new associations while not sacrificing the locality, we deliberately set $\lambda=5\times 10^{3}$ for COMEBA-HK and keep it as the optimized value for MEMIT, which are $2\times 10^{4}$ and $1.5\times 10^{4}$ for GPT2-XL and GPT-J respectively.
The evaluation results are shown in Table 1. COMEBA-HK outperforms all baselines except the reliability metric for GPT-J. Specifically, despite the relatively low $\lambda$, our method overwhelms other baselines in reliability and generality metrics for GPT2-XL meanwhile maintaining a better locality. This indicates that lowering $\lambda$ or, in other words, increasing the weight of the new associations does not sacrifice the locality in COMEBA-HK, which demonstrates the all-sided improvement of our method. The improvement margin is less for GPT-J compared with those for GPT2-XL, yet our method still ranks first for most of the metrics in both datasets. Interestingly, the reliability in the COUNTERFACT dataset for GPT-J is lower than that of MEMIT and even lower than the performance of GPT2-XL. We believe this may be one side-effect of descending $\lambda$, which inevitably gives rise to the increase of weight change magnitude at each batch editing step and consequently distorts previously updated associations more severely (sec.4.5). However, this phenomenon does not apply to consecutive editing cases, which is shown in the next section.
4.3 Evaluation on Consecutive Batch Editing
We evaluate our method’s capability on 1k samples from both datasets for consecutive batch editing, i.e., there is no roll-back. The evaluation is conducted after the end of the whole consecutive batch editing process. We set $\lambda$ to $15,000$ as the scenario now is consecutive batch editing and set the batch size to 10, that is, 100 batch editing steps.
Results are shown in Table.2. Our method surpasses other methods in 100 consecutive steps with an even larger improvement margin for almost all the metrics compared to the single-round batch editing. This demonstrates that our method does not depend on simple regurgitation of the editing samples nor rely heavily on the trade-offs of lowering the balancing factor $\lambda$ to increase the reliability and locality.
We extend the consecutive steps to 1k to explore the limit of our method. Results can be found in Fig.7. Surprisingly, the locality experiences a great fall from 1k to 2k steps but remains steady from 200 to 1k editing steps, which proves that the hook layer stably obstructs the out-scope samples. Reliability and generality consistently fall as the consecutive steps grow, indicating that there is still room for improvement in this field.
4.4 Validation of Local Editing Scope
Figure 3: Difference in z-score between updated keys and others.
Given an updated hook layer with the weight $W_{h}$, the original model weight $W_{0}$, an updated key $k_{i}$, and an out-of-scope key $k_{j}$, we conduct experiments to verify whether $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ holds. We select 100 samples from the COUNTERFACT dataset to edit GPT2-XL using COMEBA-HK, then apply the edited model to these 100 samples and record the z-score to the L2-norm of the difference vector between the last hook layer response and original model response of the updated key (the last subject token), namely, z-score of $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$. Both reliability and generality prompts are included for comprehensiveness.
The result is shown in Fig.3. Almost all the z-scores of the responses from updated keys exhibit a great margin from the mean value, with the lowest around 1.5 in reliability samples and 2 in generality samples. This not only validates our editing scope identification (section 3.1.3) technique but also cleans a possible doubt about the undesired block of generality instances in the hook layer. In addition, the discriminative z-score could also ensure the preciseness of the swapping, which can improve the locality.
4.5 Detailed Analysis and Discussions
We further conduct ablation studies and detailed analyses of different settings of the proposed method. We investigate cases with different sample sizes that correspond to different numbers of consecutive steps for each experiment. Note that the evaluation is conducted after the corresponding entire consecutive steps are finished. For example, in Fig.4, we conduct experiments for sample sizes 200, 400, 600, 800, and 1000, so the evaluation is triggered right after the first 200, 400, 600, etc, samples are edited to the model. Unless specified, the batch size is selected to be 10.
Ablation Study of Update Mechanism and Hook Layers
Figure 4: Ablation study for update mechanism and hook layers.
The effectiveness of the derived consecutive updating mechanism and the hook layers are discussed in this part. We run three cases using GPT2-XL, namely, MEMIT (no consecutive updating mechanism, no hook layers), COMEBA without hook (COMEBA w/o HK), and COMEBA-HK for consecutive batch editing on 1k samples from both ZsRE and COUNTERFACT datasets.
The results are demonstrated in Fig.4. In almost all metrics of the two datasets except the generality of COUNTERFACT, the COMEBA w/o the hook performs better than the vanilla MEMIT, and the margin tends to increase as the consecutive steps ascend. This certifies the effectiveness of our derived consecutive updating mechanism in consecutive batch editing scenarios. For the ZsRE dataset, the method with hook layers considerably outperforms the one without hook in the locality without sacrificing reliability and generality. This verifies that the hook layer can efficiently and accurately block the out-scope instances from the input without fraudulently missing in-scope instances. For the COUNTERFACT dataset, the reliability of COMEBA-HK is consistently higher than the other two, and the generality surpasses that of MEMIT after 80 editing steps. It seems that the hook layer causes some side effects in the locality of COUNTERFACT, but this circumstance is not found in the ZsRE dataset. It is worth noting that COMEBA-HK shows the most stable performance as the number of consecutive editing steps grows, demonstrating the great potential of our method for consecutive editing.
Effect of the Number of Editing Layers
Figure 5: Performance comparisons on different number of editing layers. Layers are selected from the critical path identified in (Meng et al., 2023).
To investigate the necessity of applying the hook layer onto multiple transformer layers, we conduct the consecutive batch editing experiment on the ZsRE dataset for GPT2-XL (Fig.5). As the effect of choosing different layers has already been studied in (Meng et al., 2023), we focus only on the effect of the number of layers. We selected the last one, three, and all number of layers from the critical path identified in (Meng et al., 2023; Yao et al., 2023), respectively.
As shown in Fig.5, the one-layer case significantly underperforms the other two cases in most of the metrics for the two datasets, which directly certifies the necessity of the expansion. In ZsRE, the difference between the performance for one layer and multiple layers tends to enlarge in reliability and generality as the consecutive editing steps increase. This may serve as evidence of our assumption in section 3.2, which mentions that the latter hook layer may capture the in-scope instances missed in former hook layers. Additionally, the all-layer case has slightly better generality than the three-layer case, and they do not show a remarkable difference in locality and reliability. A similar situation could be found in the reliability and generality of the COUNTERFACT with an interesting exception in the locality, where an adverse performance order of the cases is shown. Nevertheless, the margin of the locality fall is not that manifest in contrast with the advancement in reliability and generality.
Effect of the Balance Factor $\lambda$
Figure 6: Performance comparisons on different values of $\lambda$ on 1k samples.
Figure 7: Performance comparisons on different values of $\lambda$ on 10k samples.
We test the effect of different $\lambda$, the balance factor between pre-training and newly updated associations. We first evaluate the COMEBA-HK with different $\lambda$ on 1k samples from ZsRE (Fig.6). It seems that a small value of $\lambda=1,000$ would cause significant damage to all three metrics, especially the reliability and generality, since they experience a great drop as the consecutive steps increase. This may result from the overly high magnitude of the weight change caused by the low value of $\lambda$, which severely distorts the previously updated associations. Meanwhile, a too-high value of $\lambda=20,000$ seems also not to be a good choice as well, which gives rise to an overly small magnitude of the weight change so that it fails to deliver the new optimized values for keys. The cases of $\lambda=5000,10000,15000$ do not show an apparent difference, so we extend further the sample size to 10k (Fig.7).
Extended results show that 5000 is not a good choice for large-consecutive editing steps, though it performs no worse than the other two in early 1k samples. The case of $\lambda=15,000$ ranks first in reliability and generality. Although it has worse performance in locality compared to $\lambda=10,000$, the margin between them gradually narrows as the consecutive steps rise. Overall, we believe 15,000 would be a reasonable selection.
5 Related Work
Recent years have witnessed prosperous development in the field of model editing. According to (Yao et al., 2023), the proposed methods so far can be generally classified into two groups, i.e., those directly modify the model’s weight, and others that do not.
The methods that do not directly alter the model weights generally follow two directions: they either employ an external memory or introduce additional adjustable parameters. Methods like T-Patcher (Huang et al., 2023) and CaliNET (Dong et al., 2022) apply new neurons that are responsible for specific mistakes in the last layer of the FFN model. Grace (Hartvigsen et al., 2022) introduces a timely adjusted code book to edit the model’s behavior. Another group of methods like (Mitchell et al., 2022b) integrates an explicit external memory as edit descriptors to help editing scope recognition.
Those directly altering the model’s weight either train a hyper-network to predict the change required by the edits (Mitchell et al., 2022a; De Cao et al., 2021) or first locate corresponding parameters that are responsible for specific knowledge and then edit the located parameters (Meng et al., 2023, 2022; Li et al., 2023; Dai et al., 2022).
6 Conclusion
In this work, we introduce a novel model editing method, COMEBA-HK, which advocates the more practical consecutive batch model editing. COMEBA-HK uses an expanded editing mechanism to support consecutive editing and newly proposed hook layers to identify the editing scope. Compared to existing model editing methods, COMEBA-HK does not require large external memory nor extra training for meta-networks or classifiers. It only needs an extra tiny amount of memory to store the edited hook layer. Comprehensive experiments are conducted to verify the effectiveness of the method over both single-round and consecutive batch editing. We also investigate the validity of the editing scope recognition in hook layers and the influence of each parameter in our method.
Impact Statements
The goal of our method is to conveniently add or fix knowledge in the LLMs without expensive retraining so that time and resources can be saved. However, we caution readers that our method should not be used to edit any malicious or offensive knowledge, including but not limited to, political authoritative facts, rumors, ethical discrimination statements, etc. It is worth reminding that the safe and responsible application of our method is really important. None of offensive, toxic, or malicious editing should be allowed.
Acknowledgements
Do not include acknowledgements in the initial version of the paper submitted for blind review.
If a paper is accepted, the final camera-ready version can (and probably should) include acknowledgements. In this case, please place such acknowledgements in an unnumbered section at the end of the paper. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix
Figure 8: A sample from ZsRE dataset.
Figure 9: A sample from COUNTERFACT dataset.
Appendix A Dataset and Implementation Details
Datasets
Fig.8 shows an example from the ZsRE dataset. Each record in ZsRE contains the subject string, the factual prompt used for testing reliability, the rephrase prompt used for generality evaluation, and the locality prompt used for evaluating the locality. Note that what the locality demands the post-edit model does is not to predict the ground truth but whatever the initial base model predicts. Similarly, the fact, rephrase, and locality prompts of each record in COUNTERFACT are applied to the evaluation of the three metrics respectively (Fig.9).
Fine-tuning
We generally follow the procedure conducted in (Meng et al., 2023, 2022). We fine-tune the $mlp_{proj}$ parameter from layer 0 for GPT2-XL and layer 21 for GPT-J since they are found to have the optimal performance. For both models, we conduct 25 optimization steps using Adam optimizer (Kingma & Ba, 2015) and use learning rate $5e^{-4}$. All other parameters of both models follow default settings.
MEND
MEND (Mitchell et al., 2022a) conducts the editing by manipulating the language models’ gradient. It trains a meta-network that employs a rank-1 decomposition of the model gradients and predicts a new rank-1 update to the corresponding model weights. In this work, we train two meta-networks using corresponding training split in the ZsRE and COUNTERFACT datasets for GPT2-XL following the default hyperparameter settings. Due to the large required computation resource for training GPT-J (6B) meta-network, we do not perform training for GPT-J. Instead, we directly download the available MEND weights for GPT-J provided by (Meng et al., 2022) and use that for the evaluation of MEND.
SERAC
Mitchell et al. (2022b) designed a memory-augmented editing method, which requires an external cache to store explicit editing cases. It also adopts a scope classifier that determines whether an input sample falls in the editing scope and a small counterfactual model for editing the in-scope cases. For both GPT2-XL and GPT-J, we train two separate models for the two datasets respectively. Following the original paper, we apply distilbert-base-cased (Sanh et al., 2019) for the scope classifier across all models. For the small counterfactual model, we employ GPT2 for GPT2-XL and gpt-j-tiny-random [^2] https://huggingface.co/anton-l/gpt-j-tiny-random for GPT-J. All hyperparameters follow default settings.
MEMIT
MEMIT (Meng et al., 2023) treats the feedforward layer of transform as a linear associative memory and uses a minimum square error optimization to add new key-value associations to layer weights. We follow the original paper to edit the layers that are in the identified critical path and set the balance factor $\lambda$ to the optimal value found in the original work. Other parameters for the two models are all set based on configurations in (Meng et al., 2023, 2022).
COMEBA-HK
COMEBA-HK expands the update mechanism in MEMIT to consecutive cases and applies hook layers to separate the weight change from the original model layer. For both models, we set $\lambda=15,000$, $\alpha_{z}=2.2$. Unless specified, we evaluate our method on full critical path layers identified in (Meng et al., 2023). We employ the same procedure in MEMIT (Meng et al., 2023) to compute the updating keys and the target values, except that the most recently updated model during the process of consecutive editing is applied for relevant computations. We applied ”torch.float16” for the GPT-J model for all experiments.
Appendix B Analysis of Hyperparameters
Effect of the Initial Threshold $\alpha_{z}$
Figure 10: Performance comparisons on different $\alpha_{z}$.
$\alpha_{z}$ is the initialization value of $\alpha$ used in the identification of local editing scope (section 3.1.3). We study its influence in this part. According to Fig.10, although the $\alpha_{z}=1$ case ranks the highest in the first 60 editing steps in generality, it consistently performs the worst in locality, indicating that it fails to intercept many out-scope inputs. This implies that 1 may be too low for the initialization. Other cases do not show noticeable differences in the three metrics since $\alpha_{z}$ is just the initial value and $\alpha$ is determined dynamically. It seems that overly low $\alpha_{z}$ would damage the hook layer’s capacity to discriminate in-scope and out-scope samples. Considering the unpredictable consecutive steps that our method may be applied, we select a relatively low value between 2 and 3, namely, $\alpha_{z}=2.2$.
To verify the significance of the dynamical determination process, we also test the fix $\alpha$ case. We chose the value of 3 as it is the standard threshold used in standardization to detect outliers. The results reveal a dramatic decline in reliability and generality and perfect fulfillment in the locality, indicating the fact that almost all of the instances are indiscriminately obstructed by the hook layers regardless of the editing scope. Besides, it is practically unrealistic to choose an optimal fixed $\alpha$ before the editing process. Therefore, we believe it is more reasonable to decide $\alpha$ dynamically.
Effect of Editing Batch Size
Figure 11: Performance comparisons on different editing batch sizes on 1k samples of ZsRE.
Does the batch size parameter affect the performance of our method? We investigate the effect of batch size by conducting single-round editing on 1k samples from ZsRE. We tested batch sizes 10, 100, and 1000 (Fig.11).
The results showcase that while a decrease in the three metrics exists as the batch size rises, the margin could be negligible, denoting that our method possesses the mass-editing capacity.